ambiee:

What is the result of 4 divided 8/12? 3 6 1/6 3/8

ambiee:

1 attachment
Zepdrix:

\[\rm\large 4\div \frac8{12}\quad=\quad 4\times\frac{12}8\quad=\quad \frac41\times\frac{12}8\quad=\quad \frac{48}8\]

Zepdrix:

Oh you have a picture for this problem? :o hmm

ambiee:

so would it be 1/6

Zepdrix:

I'm not sure how that picture relates to this problem :d hmm

Zepdrix:

48/8 does not equal 1/6... no :u

ambiee:

yea if u simplify it it equals to 6/1

Zepdrix:

Good, and 6/1 is the same as 6, right? :)

ambiee:

okay i got another one

ambiee:

Which division problem does the diagram below best illustrate?

1 attachment
ambiee:

16 divide 4 =4 32 divide 4 =8 36 divide 4 =9 8 divide 2 =4

Zepdrix:

Hmm if you count the squares individually, you should get 32. And in the image those squares appear to have been "divided" into groups of 4. As a result, we end up with 8 of those groups.

ambiee:

AYEEE SMART

ambiee:

HERES ANOTHER ONE

ambiee:

Daniel wants to do 312 practice problems to prepare for an exam. If he wants to do approximately the same number of problems each day for 36 days, which best describes the number of problems that he needs to do each day?

1 attachment
ambiee:

8 problems per day for 12 of the days and 9 problems per day for 24 of the days 8 problems per day for 24 of the days and 9 problems per day for 12 of the days 8 problems per day for each of the days 9 problems per day for each of the days

Zepdrix:

312/36 = 8.66666... Or if we were to write it as a quotient and remainder, 312/36 = 8 + remainder of 24 Hmm this problem is a little trickier. So for 36 days, he can read 8 pages the first 12 days, and then for the next 24 days, he'll have to read one extra page per day (in order to deal with that remainder value). so he'll read 9 pages the other 24 days.

ambiee:

youre smart are u good at science

Hero:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Zepdrix I'm not sure how that picture relates to this problem :d hmm \(\color{#0cbb34}{\text{End of Quote}}\)